### v1

@@ 3.2.1. Independent Variables
If highlighting is enabled, instructing the AI to modify a part of the object will cause an outline to appear around the object part in question. If none of the other visualization aspects are enabled, this outline will be displayed for three seconds and then disappear at the same time that the object part is replaced with its modified version. If incremental visualization is enabled, the highlight will be displayed on the object part for exactly as long as the incremental process takes. If embodiment is enabled, the highlight will be displayed on the object part from the moment the AI avatar starts walking toward the object part until it has finished stepping back from the object. We base this concept of highlighting active elements on traditional 2D interface design for guiding user attention.
@@
If incremental visualization is enabled, instructing the AI to modify an object part will cause this part to disappear and its modified version to slowly appear slice by slice as if a 3D printer was printing it. If the embodiment is enabled, there will be a ”printing layer” placed on top of the incrementally growing object part, vaguely resembling the additive manufacturing technique of sheet lamination. The AI avatar will appear to move this layer along the printing direction through a line connecting its right hand to one of the corners of the layer.
@@
We varied all three independent variables in a repeated measures design, resulting in a total of conditions. We counterbalanced the order of all factors in a Balanced Latin Square to prevent learning effects. Each participant conducted eight modeling trials across eight different conditions. In each session, participants modified one of the four 3D objects (see 6(d)). Each object was presented to each participant twice, with the order randomized for each individual. We chose four objects as a suitable compromise, striking a balance between variety and familiarity. This approach allowed participants to engage their creativity by attempting various design approaches with the same model, while also providing a second chance to revisit and revise previous modifications – an opportunity to learn from failure that we argue is valuable in our research context.
| Overall experience | How was you experience carrying out the given task? |
@@ 3.2.2. Dependent Variables
During modeling sessions: A log of user interactions with the system was recorded. This consisted of the type of each interaction with the object and the AI as well as the time at which it was performed. Additionally, a screencast of the participant’s perspective in VR was recorded in each modeling session.
@@
The Inclusion of Other in the Self (IOS) Scale (Aron et al., 1992) is a pictorial single-item scale measuring closeness between the respondent and another person. Though the AI in our co-creative system is not a person, we are nevertheless interested in whether closeness, when applied to a co-creative AI partner, is affected by our visualization aspects. We therefore included this item in our questionnaire.
@@
In addition to the items sourced from the above-mentioned questionnaires, we included three Likert items on AI competence, contribution value, and AI creativity, amounting to 27 questions in total111 A list of all items is in the supplementary materials..
@@ 3.4. Procedure
After welcoming participants and before starting the study, participants filled out an informed consent form and completed a demographic information survey. We then provided an overview of the study procedure and system usage guidelines. Next, we instructed the participants on using the HMD and controllers and let them familiarize themselves with the VR interactions in a virtual room by testing selection and painting operations on a sphere. Once ready, participants began their first modeling session, tasked to ”be creative and work together to create a new version of this model”, emphasizing a focus on making the given object look different from its default state, but letting participants choose their own design goals. We timed a 5-minute session (visualized by a virtual clock) and automatically disabled all interactions once time ran out. Participants were invited to voice any thoughts they had during the sessions Participants then filled out questionnaires outside the VR environment before repeating this process for a total of eight conditions. After completing all sessions, we conducted a semi-structured interview with each participant. In total, the study lasted around 90 minutes. All participants were given cake as compensation for their time.
@@ 3.5. Participants
We recruited 16 participants (7 identified as female, 9 as male, 0 non-binary). Ages ranged from 20 to 30, with a mean age of 24 (). Out of all the participants, five had no prior experience with VR, while nine had tried it before and two had used it more often. Nine participants had no experience with 3D modeling, while three had tried it before, and four had done it more often. Most participants had interacted with AI in the form of Large Language Models before, while half of participants had used AI for image generation before, and only three participants had experience with AI for 3D object generation.
@@ 4. Results
In this section, we report the quantitative and qualitative results of our empirical study.
@@ 4.1.1. Number of Engagements
We analyzed the number of engagements as a measure of the participants’ interest in contributing to the model (see Figure 4). We defined the number of engagements as the total number of selections, modifications, and undo-actions that a user performed. We found results ranging from , (with highlighting and incremental visualization without embodiment) to , (without any ai contribution visualization).
@@ 4.2.1. Perception of AI
We assess the perception of the AI in general by analyzing the questions on predictiability, support, efficiency, and competence. None of the visualization aspects had a significant effect on predictability, efficiency, or competence. However, in conditions where embodiment was enabled, participants rated the system as significantly more supportive (, ).
@@ 4.2.2. Perception of AI output
We assess the perception of AI output by analyzing the questions on unexpectedness, unusualness, value, and creativity. There were no significant differences in ratings for unusualness, value, or creativity for any of the visualization aspects. However, in conditions where incremental visualization was enabled, participants rated the finished 3D model as significantly more unexpected (, ).
@@ 4.2.3. Collaborative experience
We assess the collaborative experience by analyzing the questions on partnership, communication, alignment, and agency. There were no significant differences in ratings for alignment or agency. However, when embodiment was enabled, participants rated communication significantly higher (, ). Additionally, participants rated partnership lower when highlighting was enabled (, ) and higher when embodiment was enabled (, ). We also found a significant interaction of highlighting and embodiment concerning this measure (, ). Post-hoc tests do not show significant differences in ratings between the individual variable combinations.
@@ 4.2.4. Appeal
We assess the overall system appeal by analyzing the questions on excitement, movitation, affinity, and enjoyment. There were no significant differences in ratings for motivation and affinity. However, in conditions where highlighting was enabled, participants enjoyed using the system significantly less (, ). Additionally, we found a significant interaction of highlighting and embodiment concerning the measure of excitement (, ). Post-hoc tests do not show significant differences in ratings between the individual variable combinations.
@@ 4.2.5. Behavioral engagement
We assess the behavioral engagement by analyzing the questions from the Behavioral Engagement Scale. Only two questions on showed significant differences between the enablement of any visualization aspects: The extent to which the AI affected what the participant did was larger when incremental visualization was enabled (, ), and the amount of attention participants paid to the AI was larger when embodiment was enabled (, ). We also found a three-way interaction of all visualization aspects concerning the latter variable (, ); however, post-hoc tests do not show significant differences in ratings between the individual variable combinations.
@@ 4.2.6. Closeness, satisfaction, and contribution
We assess these factors by analyzing the questions on satisfaction, contribution, and closeness. There were no significant differences in closeness to the AI as measured by the IOS Scale. However, in conditions where highlighting was enabled, participants were significantly less satisfied with the finished 3D model (, ). In conditions where embodiment was enabled, participants attributed significantly more of the contribution to the finished 3D model to the AI (, ).
@@ 4.3.1. Data Analysis
We collected qualitative data in a final semi-structured interview. Our data analysis procedure is based on Braun and Clarke’s (Braun and Clarke, 2006) step-by-step approach of thematic analysis. While the analysis was conducted by two of the authors, the results and intermediate steps were discussed among all authors. The analysis was performed based on the transcripts of the audiorecorded interviews. The interview responses were a mix of English and German. For the purpose of presenting the data here, all quotes are translated to English by the authors, who are fluent in English and in German. The two authors first familiarzed themselves with the data by reading through the transcripts. We used Delvetool 222Delvetool for qualitative data analysis https://app.delvetool.com/, last accessed: August 30, 2024 for coding the data. In the first analysis step both authors each read the same subset of transcripts (8 transcripts/50%) separately and created initial codes inductively. They then discussed the codes and started forming themes by grouping the codes using an online mind map tool 333Miro tool for online mind maps: https://miro.com/, last accessed: August 30, 2024. After the initial theme generation, all authors discussed the themes before one of the authors coded the remaining transcripts using the updated codes. During the second coding stage, some new codes arose and were merged with the existing themes. Finally, the two authors met again to discuss the final codes, and update and name the themes.
4.3.2. Codes and Themes
Coding the first half of the transcripts produced 103 codes. They were then grouped into 21 initial summary codes. Before coding the second half of the transcripts, we discussed and summarised the codes into 73 unique codes. Next, the second half of the transcript were coded, during which we added 15 codes to the coding set and merged them with the existing ones into 20 summary codes. We then formed the following four themes based on the summary codes444The mind map of the codes is available in the supplementary materials.
4.3.3. Task Division and Control
Participants experienced varying levels of control over the AI, ranging along a spectrum from dependency to independency and different hierarchical levels. One key aspect discussed by participants was how to divide tasks between themselves and the AI, particularly concerning the level of control each would have (or desire to have) over the task. Their comments revealed a relationship axis, involving both a degree of dependence/independence and, in case of dependence, a hierachry of interaction. The participants’ comments were quite diverse, showing no common agreement on the type of control they felt over the AI. Some participants (P13, P15) mentioned that they did not feel they were working together with the AI. Others noted that they felt like they were co-working with the AI but not actually collaborating (P8): ” It felt more like … when you’re [..] doing exercise sheets, and you have a partner who never talks to you and just hands something in without you even noticing.”. Some participants also stated that they simply ignored the input from the AI (P15). Although the AI’s input was indeed predetermined, some users felt in control of its actions and described a hierarchical relation (P1). They viewed the AI more as a tool rather than as a collaborator or co-worker. In this context, participants felt that the AI should support the user. The third relationship we observed was that of a partner, where participants felt they were pursuing a common goal with the AI. While many participants did not note any particular influences of the embodiment on the relationship, P5 explained that they felt the feeling of collaboration was more relevant to them than working with an embodied AI. Finally, some participants described feeling that the AI was in control and that their decisions were overruled by it (P2, P8).
Participants also wished for different types of relationships when asked for improvements or desired changes of the interaction. Some participants suggested a form of task sharing with the AI, where both would choose from a set of options (P4). Other participants desired a more hierarchical way of control, where they wanted to give commands to the AI (P1, P2), while others were okay with the AI working on its own, as long as the user could make the final decision about accepting or rejecting its decisions (P2): So, [the AI] should be able to try out things, but in any case it should not override my decisions. These wishes are closely related to the desire for a better communication between AI and user (P3, P6). For example, the AI should indicate whether they understood the user’s request and whether the requested action was possible (P3). P2 also indicated that they would like to receive feedback from the AI about its perception of the task.
In addition to spoken or verbal feedback, participants indicated that they would like to interact with the AI through different modalities and meta level instructions, for example, by pointing on things or by making general style requests that are then performed by the AI (P2: ”for example I can write: I want to make a goth house, and then the AI automatically suggests goth windows and so on.).
4.3.4. AI as Enabler of New Experiences
Interacting with the AI sparked creativity among participants and enabled new experiences. We observed that some participants felt the AI enabled participants to have more creative experiences. For example, participants liked the AI’s suggested modifications (P1). P8 liked that these suggestions guided them towards a satisfying end result and P7 appreciated that the AI suggestions led to new inputs: […] so when they [the AI] conjured something new, then of course one got new input. P2 also mentioned that they felt inspired by the AI: […] the things from the AI also inspired me during the process. Some participants even mentioned that they felt the AI helped them to be more creative and open (P7, P8), stating that they were happy with the result although they, for example, chose a color scheme that they would not have gone for themselves.
4.3.5. Emotional Experience with AI
Participants had both positive and negative experiences with the AI. In general, participants found the experience interesting (P2, P6, P7). Their opinions and experiences with the AI were both positive and negative. On the positive side, participants felt that the AI created funny results and a funny experience in general (P3, P4). On the negative side, participants mentioned that they were confused about the AI and its created outcome (P2). They also did not like the slow response time or having to wait for the AI to finish their modifications (P4, P5). Finally, some participants felt surprised (P3, P4) by the suggestions or distracted by the AI output (P1, P3).
4.3.6. Experiences with Embodied AI
The embodied AI evoked a variety of emotions. The participants experienced a quite diverse set of emotions when interacting with the embodied AI. Some found it annoying, because they had to wait for the embodied AI to finish the animation (P5), and disliked it being slow (P3, P4, P5, P7). While some participants liked to watch the AI, others felt it was distracting (P3). However, participants also felt curiosity when working with the embodied AI (P8).
Despite these negative feelings, participants also expressed positive emotions towards the AI. For example, they perceived it to be helpful (P1), cute (P2), or funny (P2). Participants also made comments about the appearance of the embodied AI. Some participants found it uncanny and weird (P3). Some participants specifically criticized the avatar’s permanent smiling expression, explaining that ”smiling for no reason is creepy” (P11) and mentioned that they did not prefer the AI to be human-like (P7). Indeed, several participants mentioned that they would like to be able to customize the appearance of the embodied AI (P3). P5 found it even scary when seeing the AI for the first time: ”Yes, I would say that during the iterations I got used to it, but all in all, I was quite scared when I saw it the first time”. However, we also observed positive interaction effects. P1 stated that ”changes felt more tangible when done by the embodied AI”. P4 felt that social conventions applied to the interaction with the AI: ”it was nicer to wait, when at least there was this person who made something”, and P7 stated that the embodied AI helped to not feel alone. Overall, some participants felt a connection to the embodied AI. P7 even said that they felt they missed the embodied AI a bit when it disappeared for some of the iterations. Finally, while some participants did not like the slow version of the continuous interaction mode, they felt it was more acceptable when the AI was embodied versus when it was not.
4.4. Participants’ Approach and Task Limitations
This section summarizes how participants generally approached the task. It also describes limitations that were introduced by the study task but are not related to the human-AI interaction.
Participants often criticized the fact that only a limited set of modifications was possible for the given models and that possible modifications did not vary when a model appeared for a second time. Some participants hoped for the ability to influence the AI’s modifications through their behavior. P11 voiced the name of the object that he was hoping to appear in a modification spot, while some participants tried to elicit specific modifications using the colors they painted with, and P08 even wrote a modification prompt onto the model itself.
Participants were also bothered by the fact that once they had undone a modification, it could not be re-instantiated. They moreover criticized several aspects of the painting mechanism, such as the limited amount of colors, the lack of an eraser, and the issue that if a model had repeating textures, painting on one part of a model could make a copy of the painted stroke appear at another location on the model that used the same texture.

### v3

@@ 3.2.1. Independent Variables
If highlighting is enabled, instructing the AI to modify a part of the object will cause an outline to appear around the object part in question. If none of the other visualization aspects are enabled, this outline will be displayed for three seconds and then disappear at the same time that the object part is replaced with its modified version. If incremental visualization is enabled, the highlight will be displayed on the object part for exactly as long as the incremental process takes. If embodiment is enabled, the highlight will be displayed on the object part from the moment the AI avatar starts walking toward the object part until it has finished stepping back from the object. We base this concept of highlighting active elements on traditional 2D interface design to guide user attention.
@@
If incremental visualization is enabled, instructing the AI to modify an object part will cause this part to disappear and its modified version to slowly appear slice by slice as if a 3D printer was printing it. If the embodiment is enabled, a ”printing layer” will be placed on top of the incrementally growing object part, vaguely resembling the additive manufacturing technique of sheet lamination. The AI avatar will appear to move this layer along the printing direction through a line connecting its right hand to one of the layer’s corners.
@@
| Overall experience | How was your experience carrying out the given task? |
@@
We varied all three independent variables in a repeated measures design, resulting in a total of conditions. We counterbalanced the order of all factors in a Balanced Latin Square to prevent learning effects. Each participant conducted eight modeling trials across eight different conditions. In each session, participants modified one of the four 3D objects (see 6(d)). Each object was presented to each participant twice, with the order randomized for each individual. We chose four objects as a suitable compromise, striking a balance between variety and familiarity. This approach allowed participants to engage their creativity by attempting various design approaches with the same model while also providing a second chance to revisit and revise previous modifications – an opportunity to learn from failure that we argue is valuable in our research context.
@@ 3.2.2. Dependent Variables
During modeling sessions: A log of user interactions with the system was recorded. This consisted of the type of each interaction with the object and the AI, as well as the time at which it was performed. Additionally, a screencast of the participant’s perspective in VR was recorded in each modeling session.
@@
The Inclusion of Other in the Self (IOS) Scale (Aron et al., 1992) is a pictorial single-item scale measuring closeness between the respondent and another person. Though the AI in our co-creative system is not a person, we are nevertheless interested in whether closeness, when applied to a co-creative AI partner, is affected by our visualization aspects. We, therefore, included this item in our questionnaire.
@@
In addition to the items sourced from the above-mentioned questionnaires, we included three Likert items on AI competence, contribution value, and AI creativity, amounting to 27 questions in total111See the supplementary materials for a list of all items.
@@ 3.4. Procedure
After welcoming participants and before starting the study, participants filled out an informed consent form and completed a demographic information survey. We then provided an overview of the study procedure and system usage guidelines. Next, we instructed the participants on using the HMD and controllers and let them familiarize themselves with the VR interactions in a virtual room by testing selection and painting operations on a sphere. Once ready, participants began their first modeling session, tasked to ”be creative and work together to create a new version of this model”, emphasizing a focus on making the given object look different from its default state but letting participants choose their own design goals. We timed a 5-minute session (visualized by a virtual clock) and automatically disabled all interactions once time ran out. Participants were invited to voice any thoughts they had during the sessions. Participants then filled out questionnaires outside the VR environment before repeating this process for a total of eight conditions. After completing all sessions, we conducted a semi-structured interview with each participant. In total, the study lasted around 90 minutes. All participants were given cake as compensation for their time.
@@ 3.5. Participants
We recruited 16 participants (7 identified as female, 9 as male, and 0 as non-binary). Ages ranged from 20 to 30, with a mean age of 24 (). Out of all the participants, five had no prior experience with VR, while nine had tried it before, and two had used it more often. Nine participants had no experience with 3D modeling, three had tried it before, and four had done it more often. Most participants had interacted with AI in the form of Large Language Models before. In contrast, half of the participants had used AI for image generation before, and only three participants had experience with AI for 3D object generation.
@@ 4. Results
In this section, we report our empirical study’s quantitative and qualitative results.
@@ 4.1.1. Number of Engagements
We analyzed the number of engagements to measure the participant’s interest in contributing to the model (see Figure 4). We defined the number of engagements as the total number of selections, modifications, and undo actions that a user performed. We found results ranging from , (with highlighting and incremental visualization without embodiment) to , (without any ai contribution visualization).
@@ 4.2.1. Perception of AI
We assess the perception of AI in general by analyzing the questions on predictability, support, efficiency, and competence. None of the visualization aspects significantly affected predictability, efficiency, or competence. However, in conditions with embodiment, participants rated the system as significantly more supportive (, ).
@@ 4.2.2. Perception of AI output
We assess the perception of AI output by analyzing the questions on unexpectedness, unusualness, value, and creativity. There were no significant differences in ratings for unusualness, value, or creativity for any visualization aspects. However, in conditions with incremental visualization, participants rated the finished 3D model as significantly more unexpected (, ).
@@ 4.2.3. Collaborative experience
We assess the collaborative experience by analyzing the partnership, communication, alignment, and agency questions. There were no significant differences in ratings for alignment or agency. However, with embodiment, participants rated communication significantly higher (, ). Additionally, participants rated partnership lower when highlighting was enabled (, ) and higher when embodiment was enabled (, ). We also found a significant interaction of highlighting and embodiment concerning this measure (, ). Post-hoc tests do not show significant differences in ratings between the individual variable combinations.
@@ 4.2.4. Appeal
We assess the overall system appeal by analyzing the questions on excitement, motivation, affinity, and enjoyment. There were no significant differences in ratings for motivation and affinity. However, in conditions with highlighting, participants enjoyed using the system significantly less (, ). Additionally, we found a significant interaction of highlighting and embodiment concerning the measure of excitement (, ). Post-hoc tests do not show significant differences in ratings between the individual variable combinations.
@@ 4.2.5. Behavioral engagement
We assess behavioral engagement by analyzing the questions from the Behavioral Engagement Scale. Only two questions showed significant differences between the enablement of any visualization aspects: The extent to which the AI affected what the participant did was larger when incremental visualization was enabled (, ), and the amount of attention participants paid to the AI was more significant with embodiment (, ). We also found a three-way interaction of all visualization aspects concerning the latter variable (, ); however, post-hoc tests do not show significant differences in ratings between the individual variable combinations.
@@ 4.2.6. Closeness, satisfaction, and contribution
We assess these factors by analyzing satisfaction, contribution, and closeness questions. There were no significant differences in closeness to the AI as measured by the IOS Scale. However, in conditions with highlighting, participants were significantly less satisfied with the finished 3D model (, ). In conditions with embodiment, participants attributed significantly more of the contribution to the finished 3D model to the AI (, ).
@@ 4.3.1. Data Analysis
We collected qualitative data in a final semi-structured interview. Our data analysis procedure is based on Braun and Clarke’s (Braun and Clarke, 2006) step-by-step approach of thematic analysis. While two authors conducted the analysis, the results and intermediate steps were discussed among all authors. The analysis was performed based on the transcripts of the audio-recorded interviews. The interview responses were a mix of English and German. All quotes were translated into English by the authors, who are fluent in English and German. The two authors first familiarized themselves with the data by reading the transcripts. We used Delvetool 222Delvetool for qualitative data analysis https://app.delvetool.com/, last accessed: August 30, 2024 for coding the data. In the first analysis step, both authors read the same subset of transcripts (8 transcripts/50%) separately and created initial codes inductively. They then discussed the codes and formed themes by grouping them using an online mind map tool 333Miro tool for online mind maps: https://miro.com/, last accessed: August 30, 2024. After the initial theme generation, all authors discussed the themes before one of the authors coded the remaining transcripts using the updated codes. Some new codes were merged with the existing themes during the second coding stage. Finally, the two authors met again to discuss the final codes and update and name the themes.
@@
4.3.2. Codes and Themes
Coding the first half of the transcripts produced 103 codes. They were then grouped into 21 initial summary codes. Before coding the second half of the transcripts, we discussed and summarised the codes into 73 unique codes. Next, the second half of the transcript was coded, during which we added 15 codes to the coding set and merged them with the existing ones into 20 summary codes. We then formed the following four themes based on the summary codes444See the mind map of the codes in the supplementary materials.
4.3.3. Task Division and Control
Participants experienced varying levels of control over the AI, ranging from dependency to independence and different hierarchical levels. One key aspect discussed by participants was how to divide tasks between themselves and the AI, particularly concerning the level of control each would have (or desire to have) over the task. Their comments revealed a relationship axis involving a degree of dependence/independence and, in the case of dependence, a hierarchy of interaction. The participants’ comments were quite diverse, showing no standard agreement on the type of control they felt over the AI. Some participants (P13, P15) mentioned they did not feel they were working with the AI. Others noted that they felt like they were co-working with the AI but not collaborating (P8): ”It felt more like … when you’re [..] doing exercise sheets, and you have a partner who never talks to you and just hands something in without you even noticing.” . Some participants also stated that they ignored the input from the AI (P15). Although the AI’s input was indeed predetermined, some users felt in control of its actions and described a hierarchical relation (P1). They viewed the AI as a tool rather than a collaborator or co-worker. In this context, participants felt that the AI should support the user. The third relationship we observed was that of a partner, where participants felt they were pursuing a common goal with the AI. While many participants did not note any particular influences of embodiment on the relationship, P5 explained that they felt that collaboration was more relevant to them than working with an embodied AI. Finally, some participants described feeling that the AI was in control and that their decisions were overruled by it (P2, P8).
Participants also wished for different types of relationships when asked for improvements or desired changes in the interaction. Some participants suggested a form of task sharing with the AI, where both would choose from a set of options (P4). Other participants desired a more hierarchical way of control, where they wanted to give commands to the AI (P1, P2), while others were okay with the AI working on its own as long as the user could make the final decision about accepting or rejecting its decisions (P2): So, [the AI] should be able to try out things, but in any case, it should not override my decisions. These wishes are closely related to the desire for better communication between AI and the user (P3, P6). For example, the AI should indicate whether they understood the user’s request and whether the requested action was possible (P3). P2 also indicated they would like to receive feedback from the AI about its perception of the task.
In addition to spoken or verbal feedback, participants indicated that they would like to interact with the AI through different modalities and meta-level instructions, for example, by pointing at things or by making general style requests that are then performed by the AI (P2: ”for example I can write: I want to make a goth house, and then the AI automatically suggests goth windows and so on.).
4.3.4. AI as Enabler of New Experiences
Interacting with the AI sparked creativity among participants and enabled new experiences. We observed that some participants felt the AI enabled them to have more creative experiences. For example, participants liked the AI’s suggested modifications (P1). P8 liked that these suggestions guided them towards a satisfying result, and P7 appreciated that the AI suggestions led to new inputs: […] so when they [the AI] conjured something new, then, of course, one got new input. P2 also mentioned that they felt inspired by the AI: […] the things from the AI also inspired me during the process. Some participants even mentioned that they felt the AI helped them to be more creative and open (P7, P8), stating that they were happy with the result, although they, for example, chose a color scheme that they would not have gone for themselves.
4.3.5. Emotional Experience with AI
Participants had both positive and negative experiences with the AI. Participants generally found the experience interesting (P2, P6, P7). Their opinions and experiences with the AI were both positive and negative. On the positive side, participants felt that the AI created funny results and a funny experience in general (P3, P4). On the negative side, participants mentioned that they were confused about the AI and its created outcome (P2). They also did not like the slow response time or waiting for the AI to finish their modifications (P4, P5). Finally, some participants felt surprised (P3, P4) by the suggestions or distracted by the AI output (P1, P3).
4.3.6. Experiences with Embodied AI
The embodied AI evoked a variety of emotions. The participants experienced quite diverse emotions when interacting with the embodied AI. Some found it annoying because they had to wait for the embodied AI to finish the animation (P5) and disliked it being slow (P3, P4, P5, P7). While some participants liked to watch the AI, others felt it was distracting (P3). However, participants also felt curiosity when working with the embodied AI (P8).
Despite these negative feelings, participants also expressed positive emotions towards the AI. For example, they perceived it to be helpful (P1), cute (P2), or funny (P2). Participants also made comments about the appearance of the embodied AI. Some participants found it uncanny and weird (P3). Some participants specifically criticized the avatar’s permanent smiling expression, explaining that ”smiling for no reason is creepy” (P11) and mentioned that they did not prefer the AI to be human-like (P7). Indeed, several participants noted that they would like to be able to customize the appearance of the embodied AI (P3). P5 found it even scary when seeing the AI for the first time: ”Yes, I would say that during the iterations, I got used to it, but all in all, I was quite scared when I saw it the first time”. However, we also observed positive interaction effects. P1 stated that ”changes felt more tangible when done by the embodied AI”. P4 felt that social conventions applied to the interaction with the AI: ”it was nicer to wait when at least there was this person who made something”, and P7 stated that the embodied AI helped not to feel alone. Overall, some participants felt a connection to the embodied AI. P7 even said they missed the embodied AI when it disappeared for some of the iterations. Finally, while some participants did not like the slow version of the continuous interaction mode, they felt it was more acceptable when the AI was embodied versus when it was not.
4.4. Participants’ Approach and Task Limitations
This section summarizes how participants generally approached the task. It also describes the limitations introduced by the study task unrelated to the human-AI interaction.
Participants often criticized the fact that only limited modifications were possible for the given models and that potential changes did not vary when a model appeared for a second time. Some participants hoped to be able to influence the AI’s modifications through their behavior. P11 voiced the name of the object that he was hoping to appear in a modification spot. At the same time, some participants tried to elicit specific modifications using the colors they painted with, and P8 even wrote a modification prompt on the object’s surface itself.
Participants were also bothered that it could not be re-instantiated once they had undone a modification. They moreover criticized several aspects of the painting mechanism, such as the limited amount of colors, the lack of an eraser, and the issue that if a model had repeating textures, painting on one part of a model could make a copy of the painted stroke appear at another location on the model that used the same texture.